Operations on Functions

Given k(x)= √x^2-8 and p(x)=2x^2

What is k ∘ p?


Given h(x)=2000-0.32x, k(x)= √x^2-8 and m(x)=5+7x^3

What is h ∘ k ∘ m?


Thank you so much!


1
Expert's answer
2021-03-25T14:02:44-0400

1.


"k\circ p=\sqrt{(2x^2)^2-8}=\sqrt{4x^4-8}"

"k\circ p=\sqrt{4x^4-8}"

Or

"k\circ p=2\sqrt{x^2-2}"

2.


"h\circ k\circ m=2000-0.32\sqrt{(5+7x^3)^2-8}"